1mole of NO2 and 2 moles of CO are enclosed in a one litre vessel to attain the following equilibrium NO2 +CO-----NO + CO2. It was estimated that at the equilibrium, 25% of initial amount of CO is consumed. The equilibrium constant Kp is

Grade:

1 Answers

Arun
6 years ago
Since 1 mole of NO2 reacts with 1 mole of CO ,Thus 25% of CO i.e 1/2 moles out of 2 moles have reacted with 1/2 moles of NO2
 
Kc=[(1/2)(1/2)]/[(1/2)(3/2)] =1/3

Kp=Kc(RT)Δn

Δn = change in no. of moles =no. of moles of product- no. of moles of reactant
Δn = (1+1)-(1+1)=0

thus 
Kp =Kc =1/3